OpenStudy (australopithecus):

Find all complex numbers such that: z^4 = -1 How do I do this? Do I take the forth root of -1?

OpenStudy (anonymous):

take the forth root of -1 or write -1 as \(i^2\)\[z^4-i^2=0\]\[(z^2-i)(z^2+i)=0\]\[z=\pm\sqrt{i}\]\[z=\pm\sqrt{-i}\]

OpenStudy (anonymous):

second is\[z=\pm i\sqrt{i}\]actually

OpenStudy (anonymous):

and finally use the fact that\[i=\exp(\frac{\pi}{2} i)\]

OpenStudy (anonymous):

makes sense?

OpenStudy (australopithecus):

I think so I just find the roots, that shouldn't be too hard I just need to consult my notes from previous class so, z^(4) = -1 I would just write as z = -1^(1/4) and then just find all the roots by taking 2pi intervals of the first root I obtain, and check to see if the solutions are unique

OpenStudy (anonymous):

right :)